OpenStudy (anonymous):

How do you solve X=2y-z/4 to solve for Z?

OpenStudy (anonymous):

do u have more inf

OpenStudy (anonymous):

No not really, im just learning this and my teacher is no use

OpenStudy (anonymous):

i guess u have to make the z in on side and the rest to the other side by multiply ...

OpenStudy (anonymous):

x = 2y - z/4 (for z) first lets multiply the entire equation by 4 to get rid of the fraction 4x = 8y - z (now subtract 8y from both sides) 4x - 8y = -z (now multiply the equation by -1 to make x positive) -4x + 8y = z

OpenStudy (anonymous):

oops...typo...multiply by -1 to make z positive